OpenStudy (anonymous):

Factor, the simplify. sin^3(x)-sin^2(x)-sinx+1

OpenStudy (gorv):

\[\sin^2x(sinx-1)-1(sinx-1)\]

OpenStudy (gorv):

take sinx-1 out

OpenStudy (anonymous):

and can that be simplified further?

OpenStudy (cwrw238):

= (sinx - 1)(sin^2 x - 1) - now you can factor the second parentheses

OpenStudy (cwrw238):

- its the difference of 2 squares

OpenStudy (anonymous):

the second parentheses would be sin(x) +1 * sin(x) -1

OpenStudy (anonymous):

Yes : So, finally you will get like: \[(\sin(x) - 1)(\sin(x) + 1) (\sin(x)-1) \implies (\sin(x)-1)^2(\sin(x)+1)\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

check the attachment